OpenStudy (anonymous):

A car is traveling at a speed of 90 kilometers per hour. What is the car's speed in meters per second? How many meters will the car travel in 10 seconds? Do not round your answers.

OpenStudy (anonymous):

(90km/h)(1000m/1 km)(1 h/3600 s) = 25m/s (25m/s)(10s) = 250m

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No Problem

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

A parachutist's speed during a free fall reaches 135 miles per hour. What is this speed in feet per second? At this speed, how many feet will the parachutist fall during 10 seconds of free fall? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers.

OpenStudy (anonymous):

do you know how to do this one?

OpenStudy (anonymous):

(135mi/h)(5280ft/1mi)(1h/3600s) = 198 ft/s (198 ft/s)(10s) = 1980 ft